OpenStudy (anonymous):

Soybean meal is 16% protien, cornmeal is 8% protien. How many pounds of each should be mixed together in order to get 320-lb mixture that is 14% protien?

OpenStudy (anonymous):

i could have sworn i just did this problem. matter of fact i am sure of it.

OpenStudy (anonymous):

good then you can tell me how to solve it lol

OpenStudy (anonymous):

be happy to write it again in you missed it.

OpenStudy (anonymous):

sure. in english or equation? or both?

OpenStudy (anonymous):

equation so I can see how it is solved

OpenStudy (anonymous):

ok you want to end up with 14% of 320=44.8 pounds of protein. 16% will come from soybeans so if we call x the amount of soybeans you will get .16 x the rest will come from cornmeal and since you will have 320 pounds total if you have 0 amount of soybeans you will have 320 - x amount of cornmeal, 8 % of which is \[.08\times(320-x)\]

OpenStudy (anonymous):

so you get the equation \[.16x+.08(320-x)=.14\times 320=44.8\]

OpenStudy (anonymous):

now we solve for x: \[.16x+.08(320-x)=44.8\] \[.16x+25.6-.08x=44.8\] \[.08x+25.6=44.8\] \[.08x=44.8-25.6=19.2\] \[x=\frac{19.2}{.08}=240\]

OpenStudy (anonymous):

and of course that makes the amount of cornmeal 80

OpenStudy (anonymous):

so it would be 44.8 ponds of soybean and 19.2 pounds of cornmeal is that right?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

solving for x gives x = 240

OpenStudy (anonymous):

240 pounds of soybeans and therefore 80 pounds of cornmeal

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

yes clearer now

OpenStudy (anonymous):

k